OpenStudy (anonymous):

Help me prove LHS is equal to RHS? 3. 1 + sec^2x sin^2x = sec^2x

OpenStudy (anonymous):

1 can be written as (cos^2x)(sec^2x). I hope u can see the answer now

OpenStudy (anonymous):

I don't get it... :c

OpenStudy (mertsj):

Do you know that the sec is the reciprocal of the cos?

OpenStudy (anonymous):

Hang on just a minute, let me write this all out.

OpenStudy (anonymous):

Okay. I'm good. Yes, I get the reciprocals!

OpenStudy (mertsj):

Ok. Glad you got it.

OpenStudy (anonymous):

Oh wait. You misunderstood! I don't understand the problem and answer, I was saying I'm good because I got it all written out! D:

OpenStudy (mertsj):

Replace sec^2x with 1/cos^2x

OpenStudy (anonymous):

Okay, then do I need to multiply 1/cos^2x and sin^2x?

OpenStudy (mertsj):

yes. And you will have: \[\frac{\sin ^2x}{\cos ^2x}\]

OpenStudy (mertsj):

What is that equal to?

OpenStudy (anonymous):

And that makes tangent, right?

OpenStudy (anonymous):

If that makes tangent, then I can use the identity tht says 1+tan^2x=sec^2x...Right?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

Awesome. Thank you! God bless you!

OpenStudy (mertsj):

yw. You too.